OpenStudy (anonymous):

second derivative of -3/x-3

OpenStudy (anonymous):

-3/x - 3? f'(x) = 3/x^2 - 0 = 3/x^2 f''(x) = -6/x^3

OpenStudy (anonymous):

Become a fan!

OpenStudy (anonymous):

how did you get x^2?

OpenStudy (anonymous):

Become a fan! Show your thanks and then maybe I will tell ya'. :D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-3/x = -3(x)^-1 So, to differentiate that: (-3)(-1) * (x)^(-1-1) = 3x^(-2) = 3/x^2

OpenStudy (anonymous):

my bad the question was -3/(x-3)

OpenStudy (anonymous):

Write those in! -3 / (x-3) -3(x-3)^-1 f'(x) = 3(x-3)^-2 f''(x) = -6(x-3)^-3